OpenStudy (anonymous):

Domain and range of x^2+7x+12(its all under the square root)

OpenStudy (anonymous):

\[f(x)=\sqrt{x^2 + 7x + 12} \rightarrow \text{f(x) is defined }\forall x\text{ | }x^2 + 7x + 12 > 0\]

OpenStudy (anonymous):

am new to this,that is how i meant it :)

OpenStudy (anonymous):

Right, so the function is defined so long as the stuff under the radical is greater than or equal to 0. So solve the quadratic and see where it equals 0 and where it's above/below. That will tell you what x values are ok.

OpenStudy (anonymous):

i got -4 and -3

OpenStudy (anonymous):

Sounds right.. so (x+4)(x+3) So when will this be negative?

OpenStudy (anonymous):

Remember that a product is negative only when it has an odd number of negative factors.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so my domain is -4>= x >= -3 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then what is it?

OpenStudy (anonymous):

When x is very large and negative what will that product be?

OpenStudy (anonymous):

\[(-\infty + 4)(-\infty+3) = ? \]

OpenStudy (anonymous):

If you put in a value (any value in between -3 and -4, let's say -3.5 the result will be an imaginary number. The ABSOLUTE VALUE of X must be greater than -3 OR greater than -4. That means any number in between -3 or -4 must be excluded! (-inf, -3] U [-4, +inf).

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The domain is (-inf, -3] U [-4, +inf).

OpenStudy (anonymous):

range?

OpenStudy (anonymous):

my brain is a bit fried from java programming last week

OpenStudy (anonymous):

God, I don't know what is wrong with me today. The range is (-inf, -3] U [-4, +inf). The domain is X.

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

If you have a function y=f(x), x is the domain and y is the range.

OpenStudy (anonymous):

Imagine two sets. set D={A,B,C} and set R={3,5,7}. If some function "f" associates the letter A with the number 3, then f(A) =3 and A is in the Domain of f and 3 is in the range of f. Did that make any sense to you?

OpenStudy (anonymous):

ok cool, am good in math , not really in domain and range but thanks

OpenStudy (anonymous):

I understand totally. It took me a long, long time to understand the concept of domain and range. If you have the time find " Schaum's Outline Series: Theory and Problems of Set Theory and Related Topics" by Seymour Lipschutz and read page 45. It will explain everything you never wanted to know about Range and Domain and didn't know who to ask. LOL.